Exercise 10-7 (Part Level Submission) Teal Furniture Company started construction of combination office and warehouse building for its own use at an estimated cost of $5,000,200 on January 1 2017, Teal expected to complete the building by December 31, 2017. Teal has the following debt obligations outstanding during the construction period. Construction loan-12% interest, payable semi annually, issued December 31, 2016 Short-term loan-10% interest, payable monthly, and principal payable at maturity on May 30, 2018 Long-term loan-11% interest, payable on January 1 of each year. Princ pal payable on January 1, 2021 $2,007,800 1,588,200 990,600 Your answer is correct. Assume that Teal completed the office and warehouse building on December 31, 2017, as planned at a total cost of $5,167,500, and the weighted-average amount of accumulated expenditures was $3,836,900 Compute the avoidable interest on this project. (Use interest rates rounded to 2 decimal places, e.g. 7.58% for computational purposes and round final answers to O decimal places, e.g. 5,275.) Avoidable Interest 430797 Click if you would like to Show Work for this question: Open Show Work SHOW SOLUTION Attempts: 1 of 5 used Compute the depreciation expense for the year ended December 31, 2018. Teal elected to depreciate the building on a straight-line basis and determined that the asset has a useful life of 30 years and a salvage value of $297,700. (Round answer to O decimal places, e.g. 5,275.) Depreciation Expense

Please make the answer clear and bold thank you so much for the help!
Answer #1
a) Interest on construction loan = 2007800*12% = 240936
Rate of interest on other loans:
= (1588200*10%+990600*11%)/(1588200+990600) = 10.38%
Avoidable interest = 240936+(3836900-2007800)*10.38% = 430797
b) Depreciable base = 5167500+430797297700 = 5300597
Depreciation expense for 2018 = 5300597/30 = 176687